OpenStudy (anonymous):

Just wanted to make sure im doing this right :) 2^5 x 2^6 simplify and evaluate - what do ya'll get?

OpenStudy (anonymous):

2^11= 2*2*2*2*2*2*2*2*2*2*2=2048

OpenStudy (calculusxy):

First evaluate 2^5. To do this, you have to multiply 2 five times. Expression: 2*2*2*2*2

OpenStudy (anonymous):

do you not add the powers together to get 2^11 and then evaluate from there?

OpenStudy (anonymous):

2^11=2048

OpenStudy (calculusxy):

@ndeths No you don't that way. You have to do things seperatly.

OpenStudy (anonymous):

if multiplied together you add the powers

OpenStudy (calculusxy):

See the difference yourself. Do 2^5 and then do 2^6. Add the two products up. And see it.

OpenStudy (anonymous):

@calculusxy we are not adding we are multiplying the 2 terms

OpenStudy (anonymous):

the final answer is 2^5*2^6=2^11=2048

OpenStudy (calculusxy):

Sorry about that

OpenStudy (calculusxy):

Yes you are right.

OpenStudy (anonymous):

@coolaserman that's what I got! :)

OpenStudy (anonymous):

To clarify, you can do it either way. Join them by adding the exponents: \(\large 2^5*2^6 = 2^{11} = 2048\) Or do the exponents separately: \(\large 2^5*2^6 = 32*64 =2048 They are equivalent.

OpenStudy (anonymous):

\(\large 2^5*2^6 = 32*64 =2048\) Forgot to close my tags on the second line -_-

OpenStudy (anonymous):

haha thanks man ! :)

OpenStudy (anonymous):

My pleasure. =)